katia09:

The equation of line CD is (y−3) = − 2 (x − 4). What is the slope of a line perpendicular to line CD? (4 points)

carmelle:

This equation is in point-slope form, where it is written as: \[y-y _{1}=m(x-x _{1})\] \[y−3 = − 2 (x − 4)\] m = slope Looking at this, we can see that the slope is -2. To find the slope of the line that is perpendicular to this equation, we need to find the opposite reciprocal. The opposite of -2 is 2. Finding the reciprocal is simple: \[\frac{ a }{ b }\rightarrow \frac{ b }{ a }\] \[\frac{ 2 }{ 1 }\rightarrow \frac{ 1 }{ 2 }\] This means the reciprocal of 2 is 1/2. `Therefore, the slope of the line perpendicular to line CD is 1/2`

katia09:

Thank you!!